OpenStudy (anonymous):

simplify 3/4-squareroot of 13 leave your answer in surd form

OpenStudy (anonymous):

\[3/4-sqrt{13}\]

OpenStudy (anonymous):

its already simplified?

OpenStudy (anonymous):

no.....itz not........

OpenStudy (anonymous):

ok i got -1.1488

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i mean -2.8556

OpenStudy (anonymous):

ummmm answer should be in surd form.....

OpenStudy (anonymous):

surd form?

OpenStudy (anonymous):

yeah......u know surds?

OpenStudy (anonymous):

wht type of math class?

OpenStudy (anonymous):

well surds is an irrational root of a rational number..........m in year 12........

OpenStudy (anonymous):

um would it be this?

OpenStudy (anonymous):

as in answer or equation?

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

how did u get the answer?

OpenStudy (anonymous):

is it correct, or potentially correct? answer that first lol

OpenStudy (anonymous):

no....i dont think so.........

OpenStudy (anonymous):

does surd form look like that?

OpenStudy (anonymous):

ummm yes...

OpenStudy (anonymous):

i got it from calculator

OpenStudy (anonymous):

i didnt know that the calculator could simplify that equation.....

OpenStudy (anonymous):

Are you sure it's (3/4) - 13^.5?

OpenStudy (anonymous):

i was told this calculator can convert into surd form http://www.wolframalpha.com/

OpenStudy (anonymous):

Because I'm not sure how you're simplifying 13^.5.... 13 is prime

OpenStudy (anonymous):

ummm thatz is what am tryin to simplify.....itz quite hard....

OpenStudy (anonymous):

i simplfief for u and got decim,al

OpenStudy (anonymous):

the 13

OpenStudy (campbell_st):

surds are anothe name for radicals... so the answer needs to be written usign radicals.

OpenStudy (anonymous):

So is the idea we have to combine the terms somehow?

OpenStudy (anonymous):

campbell

OpenStudy (anonymous):

is this right?

OpenStudy (campbell_st):

just a quick question about the expression is it 1. \[\frac{3}{4 - \sqrt{13}}\] or 2. \[\frac{3}{4} - \sqrt{13}\]

OpenStudy (anonymous):

the first one....

OpenStudy (anonymous):

wow u should of told me

OpenStudy (anonymous):

This is why I said "Are you sure it's (3/4) - 13^.5?" lol. I need to learn to Tex on this thing,

OpenStudy (campbell_st):

ok... so you need to rationalise the denominator multiply by 1.... in another form so its and use the difference of 2 squares this will give a rational denominator \[\frac{3}{4 - \sqrt{13}} \times \frac{4 + \sqrt{13}}{4 + \sqrt{13}}\]

OpenStudy (campbell_st):

do you think you can work it out from here @prasheelasingh

OpenStudy (anonymous):

ummm yea....thanks

OpenStudy (campbell_st):

glad to help

OpenStudy (anonymous):

ummm i need help.......

OpenStudy (campbell_st):

so do the denominator 1st its the difference of 2 squares \[(a - b)(a + b)=a^2 - b^2\] you need \[a = 4 ... and... b = \sqrt{13}\]

OpenStudy (anonymous):

ummm i got 16 -13.... ????????????

OpenStudy (campbell_st):

correct... so thats means the denominator is 3 now is there a common factor in the numerator that cancels..?

OpenStudy (anonymous):

hmmmm is the answer